# Completeness of coalition logics with seriality, independence of agents, or determinism

### v1

@@ 3.4 Upward derivability lemma
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, we have $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$ $\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
@@
    Assume $\mathsf{D}\in\mathsf{X}$. Then then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\vee\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Rule $\mathtt{Mon}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{Det}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$. We respectively consider whether $\mathsf{S}$ is in $\mathsf{X}$ and whether $\mathsf{I}$ is in $\mathsf{X}$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, $\mathbb{I}^{\prime}$ is $\mathsf{D}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{IA}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{I}\mathsf{D}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
@@
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}$ $\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.

### v2

@@ 3.6 Upward derivability lemma
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$ is $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, we have $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$ $\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
@@
    Assume $\mathsf{D}\in\mathsf{X}$. Then then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\vee\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Rule $\mathtt{Mon}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{Det}$ is an axiom of $\mathsf{MCL}+\mathsf{X}$}. We respectively consider whether $\mathsf{S}$ is in $\mathsf{X}$ and whether $\mathsf{I}$ is in $\mathsf{X}$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, $\mathbb{I}^{\prime}$ is $\mathsf{D}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{IA}$ is an axiom of $\mathsf{MCL}+\mathsf{X}$}, and $\mathbb{I}^{\prime}$ is $\mathsf{I}\mathsf{D}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
@@
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$ is $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}$ $\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.